# Quantitative Attractor Analysis of High-Capacity Kernel Logistic Regression Hopfield Networks

### v2

@@ Abstract
Kernel-based learning methods such as Kernel Logistic Regression (KLR) can dramatically increase the storage capacity of Hopfield networks, but the principles governing their performance and stability remain largely uncharacterized. This paper presents a comprehensive quantitative analysis of the attractor landscape in KLR-trained networks to establish a solid foundation for their design and application. Through extensive, statistically validated simulations, we address critical questions of generality, scalability, and robustness. Our comparative analysis reveals that KLR and Kernel Ridge Regression (KRR) exhibit similarly high storage capacities and clean attractor landscapes, suggesting this is a general property of kernel regression methods, though KRR is computationally much faster. We uncover a non-trivial, scale-dependent scaling law for the kernel width ($\gamma$), demonstrating that optimal capacity requires gamma to be scaled such that $\gamma\times N$ increases with network size $N$. This implies that larger networks necessitate more localized kernels—where each pattern’s influence is more spatially confined–to manage inter-pattern interference. Under this optimized scaling, we provide definitive evidence that the storage capacity scales linearly with network size ($P\propto N$). Furthermore, our sensitivity analysis shows that performance is remarkably robust to the choice of the regularization parameter lambda. Collectively, these findings provide a clear set of empirical principles for designing high-capacity, robust associative memories and clarify the mechanisms that enable kernel methods to overcome the classical limitations of Hopfield-type models.
@@ 1 Introduction
Nonlinear kernel methods offer a powerful means to transcend these constraints. Techniques such as Support Vector Machines (SVMs), Kernel Ridge Regression (KRR), and Kernel Logistic Regression (KLR) implicitly map data into high- or even infinite-dimensional feature spaces, where complex nonlinear relationships become linearly separable [6]. Applying this framework to associative memory learning holds the potential to dramatically increase storage capacity and enhance noise robustness.
Our previous work [7] revealed the remarkable potential of Kernel Logistic Regression (KLR) for training high-capacity Hopfield networks, demonstrating storage capacities far exceeding the classical limit. This initial discovery, however, left fundamental questions regarding the generality, scalability, and robustness of this approach unanswered. For instance, is the choice of the loss function critical? How does performance scale with network size $N$? And which hyperparameters govern the stability?
This paper aims to answer these questions through a comprehensive quantitative analysis. We systematically investigate the attractor landscape by conducting a series of rigorous, statistically validated experiments, including (1) a direct comparison with Kernel Ridge Regression (KRR), (2) a multi-scale analysis of storage capacity, and (3) a detailed hyperparameter sensitivity study.
@@
    We confirm and statistically validate that KLR-trained networks achieve significantly higher storage capacity and noise robustness, exhibiting a remarkably “clean” attractor landscape with near-zero rates of spurious fixed points.
@@ 3.2.3 Kernel Logistic Regression (KLR) Hopfield Network
$L_{i}(\bm{\alpha}_{i})=-\sum_{\mu=1}^{P}\left[t^{\mu}_{i}\log(\sigma(h_{i}(\bm{\xi}^{\mu})))+(1-t^{\mu}_{i})\log(1-\sigma(h_{i}(\bm{\xi}^{\mu})))\right]+\frac{\lambda}{2}\|\bm{\alpha}_{i}\|^{2},$ (4)
@@ 3.2.4 Kernel Ridge Regression (KRR) Hopfield Network
$L_{i}(\bm{\alpha}_{i})=\sum_{\mu=1}^{P}(\xi^{\mu}_{i}-h_{i}(\bm{\xi}^{\mu}))^{2}+\frac{\lambda}{2}\|\bm{\alpha}_{i}\|^{2}.$ (5)
@@
where $K$ is the $P\times P$ kernel (Gram) matrix $K_{\mu\nu}=K(\bm{\xi}^{\mu},\bm{\xi}^{\nu})$, $I$ is the identity matrix, and $\bm{Y}$ is the $P\times N$ matrix of stored patterns. This non-iterative solution makes the training process of KRR significantly faster than the gradient-based optimization required for KLR, a difference we quantify in our experiments.
@@ 3.4 Recall Process
The recall process in KLR-trained networks differs from that in Hebbian and LLR-based networks, as it does not rely on an explicit $N\times N$ weight matrix. Given a network state $s(t)$ at time $t$, the next state $s(t+1)$ is determined by the sign of each neuron’s activation potential relative to a threshold $\theta$:
@@ 3.5.1 Network and Learning Parameters
For the KLR learning process, we used gradient descent with a learning rate $\beta=0.1$ and number of parameter updates $M=200$. For the KRR learning process, the closed-form solution was used. The L2 regularization parameter $\lambda$ was set to a default value of $\lambda=0.01$ for most experiments, except for the sensitivity analysis. As shown in Appendix A, the learning curve for a challenging high-load condition ($N=500$, $P/N=4.0$) demonstrates that the loss substantially decreases within the initial $\sim$150 steps and approaches a plateau, confirming that 200 updates are adequate to achieve a well-converged model.
@@ 3.6.3 Quantitative Metrics
  • Fixed Point Rate: The proportion of trials that converged to any fixed point (including target, other learned, and spurious). This is equivalent to 1 - (Cycle Rate + Not Converged Rate).
@@ 4 Results
A preliminary analysis revealed that the rate of convergence to limit cycles (Cycle Rate) was negligible ($0.1\%$) across all tested parameter regimes (see Appendix C for a representative plot). This confirms that the network’s dynamics are overwhelmingly dominated by convergence to fixed points. Therefore, the following sections will focus on the analysis of these fixed point attractors and their properties.
@@ 4.1 Performance and Attractor Landscape Characteristics ($N=500$)
Figure 1: Performance and attractor landscape characteristics of the KLR-trained network ($N=500$). All curves show the mean values averaged across five master random seeds, with shaded regions representing the 95% confidence intervals. (a) Capacity Curve: Target Recall Rate as a function of storage load ($P/N$) for various initial similarities. The network maintains near-perfect recall up to $P/N=4.0$ for low-noise inputs. (b) Decomposition of Recall Errors: Recall Errors (1.0 - Target Recall Rate) are decomposed into convergence to other learned patterns and spurious attractors. Colors distinguish the error type, while line styles represent different initial similarities. The rate of spurious errors is negligible, demonstrating a remarkably clean landscape. The dominant failure mode at high load and noise is mis-recall to other learned patterns. (c) Noise Robustness: Target Recall Rate as a function of initial similarity for various storage loads ($P/N$). At low loads, recall is robust even from highly corrupted states. (d) Convergence Speed: Average number of steps to convergence for various initial similarities. Recall is exceptionally fast, typically within 1-2 steps for high-similarity inputs and well under 10 steps even for noisy, high-load conditions.
@@
To understand the nature of the recall failures observed at high loads and noise levels, we decomposed the errors into their constituent sources. The results, shown in Fig. 1(b), reveal a crucial characteristic of the KLR-trained landscape. The rate of convergence to spurious attractors (Spurious Error, green lines) is negligible, remaining close to zero across all conditions. This demonstrates that KLR learning effectively suppresses the formation of unintended stable states, a primary cause of failure in traditional models. Instead, the dominant source of error is the convergence to an incorrect but valid learned pattern (Other Learned Error, orange lines). This “mis-recall” becomes the principal failure mode as the storage load and initial noise increase, indicating that the performance limit is defined not by a corrupted landscape, but by the increasing competition between the basins of attraction of the stored memories themselves.
@@
Finally, the recall dynamics are exceptionally fast, as illustrated in Fig. 1(d). For high-similarity initial states, the network consistently converges in just one to two steps. Even for highly corrupted states in the low-similarity regime, the average number of steps to convergence remains remarkably low, typically well under 10 steps. Interestingly, for these noisy inputs, the convergence time exhibits a non-monotonic “hump” shape, peaking at intermediate storage loads. The subsequent decrease in convergence time at very high loads corresponds precisely to the regime where mis-recalls become frequent (cf. Fig. 1(b)), suggesting that the dynamics quickly settle into nearby, incorrect attractors rather than undertaking a prolonged search for the correct target. This rapid convergence is a significant practical advantage of the model.
@@ 4.3 Scaling Analysis of Storage Capacity and Dynamics
Figure 3: Scaling analysis of storage capacity and convergence dynamics under N-dependent optimal kernel widths ($\gamma_{\text{opt}}$). The analysis was performed under a moderate noise condition (Initial Similarity = 0.8) for three different network sizes: $N=100,250$, and $500$. (a) Storage Capacity Scaling: The Target Recall Rate curves for all three network sizes collapse onto a single, universal curve as a function of the storage load ($P/N$). This provides strong evidence that the storage capacity $P$ scales linearly with the network size $N$. (b) Convergence Speed Scaling: In contrast, the convergence dynamics exhibit a clear dependence on the absolute network size. For a given $P/N$ ratio in the high-load regime, larger networks converge significantly faster. This suggests that while the stability of attractors is governed by relative load, the efficiency of the recall dynamics is enhanced in higher-dimensional state spaces.
@@
During our initial investigation, a naive application of the conventional kernel scaling rule, $\gamma=1/N$, led to the counter-intuitive result that the storage capacity coefficient $P/N$ decreased for larger networks. This motivated a focused analysis of the interplay between $N$ and $\gamma$. We found that the optimal kernel width, $\gamma_{\text{opt}}$, which maximizes recall performance at high loads, exhibits a non-trivial dependence on the network size. Specifically, our empirical analysis revealed that larger networks require a more localized kernel to achieve maximum capacity, with the optimal scaling factor $c_{\text{opt}}=\gamma_{\text{opt}}*N$ increasing with $N$ (e.g., $c_{\text{opt}}\approx 2$ for $N\leq 250$ and $c_{\text{opt}}\approx 5$ for $N=500$). A detailed account of this hyperparameter optimization is provided in Section 4.4.
@@
As shown in Fig. 3(a), the Target Recall Rate curves for all three network sizes now collapse beautifully onto a single, universal curve. The performance is determined solely by the relative storage load $P/N$, irrespective of the absolute network size $N$. This result provides strong evidence that, under appropriately scaled kernel parameters, the storage capacity of the KLR-trained network follows the fundamental linear scaling law, $P\propto N$. This confirms that the high capacity coefficients reported are a general property of the model and not a finite-size effect.
@@ 5 Discussion
Our extensive quantitative analysis has provided a comprehensive picture of the attractor landscape in KLR-trained Hopfield networks. A central finding is the emergence of a remarkably “clean” landscape, characterized by a near-zero rate of spurious fixed points even under high storage loads (Fig. 1(b)). This strong suppression of spurious states is a key factor behind the model’s high fidelity. But what is the mechanism behind this suppression?
@@ 6 Conclusion
Our findings confirm that KLR training sculpts a remarkably clean and robust attractor landscape, characterized by a near-zero rate of spurious states and exceptionally fast recall dynamics. Our comparative analysis revealed that this high performance is a general feature of kernel regression methods, though KLR and KRR present a trade-off between iterative flexibility and closed-form efficiency.
@@
In summary, this work provides a clear set of empirical findings and practical design principles for constructing high-capacity, kernel-based Hopfield networks. By systematically characterizing the properties of the attractor landscape and its dependence on key parameters, we have laid a solid foundation for future theoretical work and the engineering of next-generation, brain-inspired memory systems.

### v3

@@ Abstract
Kernel-based learning methods such as Kernel Logistic Regression (KLR) can dramatically increase the storage capacity of Hopfield networks, but the principles governing their performance and stability remain largely uncharacterized. This paper presents a comprehensive quantitative analysis of the attractor landscape in KLR-trained networks to establish a solid foundation for their design and application. Through extensive, statistically validated simulations, we address critical questions of generality, scalability, and robustness. Our comparative analysis reveals that KLR and Kernel Ridge Regression (KRR) exhibit similarly high storage capacities and clean attractor landscapes under typical operating conditions, suggesting this is a general property of kernel regression methods, though KRR is computationally much faster. We uncover a non-trivial, scale-dependent scaling law for the kernel width ($\gamma$), demonstrating that optimal capacity requires $\gamma$ to be scaled such that $\gamma\times N$ increases with network size $N$. This implies that larger networks necessitate more localized kernels—where each pattern’s influence is more spatially confined—to manage inter-pattern interference. Under this optimized scaling, we provide definitive evidence that the storage capacity scales linearly with network size ($P\propto N$). Furthermore, our sensitivity analysis shows that performance is remarkably robust to the choice of the regularization parameter $\lambda$. Collectively, these findings provide a clear set of empirical principles for designing high-capacity, robust associative memories and clarify the mechanisms that enable kernel methods to overcome the classical limitations of Hopfield-type models.
@@ 1 Introduction
Nonlinear kernel methods offer a powerful means to transcend these constraints. Techniques such as Support Vector Machines, Kernel Ridge Regression (KRR), and Kernel Logistic Regression (KLR) implicitly map data into high- or even infinite-dimensional feature spaces, where complex nonlinear relationships become linearly separable [6]. Applying this framework to associative memory learning holds the potential to dramatically increase storage capacity and enhance noise robustness.
Our previous work [7] revealed the remarkable potential of KLR for training high-capacity Hopfield networks, demonstrating storage capacities far exceeding the classical limit. This initial discovery, however, left fundamental questions regarding the generality, scalability, and robustness of this approach unanswered. For instance, is the choice of the loss function critical? How does performance scale with network size $N$? And which hyperparameters govern the stability?
This paper aims to answer these questions through a comprehensive quantitative analysis. We systematically investigate the attractor landscape by conducting a series of rigorous, statistically validated experiments, including (1) a direct comparison with KRR, (2) a multi-scale analysis of storage capacity, and (3) a detailed hyperparameter sensitivity study.
@@
    We confirm and statistically validate that KLR-trained networks achieve significantly higher storage capacity and noise robustness, exhibiting a remarkably clean attractor landscape, nearly devoid of spurious states under typical noise conditions.
@@ 3.2.3 Kernel Logistic Regression (KLR) Hopfield Network
$L_{i}(\bm{\alpha}_{i})=-\sum_{\mu=1}^{P}\left[t^{\mu}_{i}\log(\sigma(h_{i}(\bm{\xi}^{\mu})))+(1-t^{\mu}_{i})\log(1-\sigma(h_{i}(\bm{\xi}^{\mu})))\right]+\frac{\lambda}{2}\bm{\alpha}_{i}^{\top}\bm{K}\bm{\alpha}_{i},$ (4)
@@ 3.2.4 Kernel Ridge Regression (KRR) Hopfield Network
$L_{i}(\bm{\alpha}_{i})=\sum_{\mu=1}^{P}(\xi^{\mu}_{i}-h_{i}(\bm{\xi}^{\mu}))^{2}+\frac{\lambda}{2}\bm{\alpha}_{i}^{\top}\bm{K}\bm{\alpha}_{i}.$ (5)
@@
where $\bm{K}$ is the $P\times P$ kernel (Gram) matrix $K_{\mu\nu}=K(\bm{\xi}^{\mu},\bm{\xi}^{\nu})$, $I$ is the identity matrix, and $\bm{Y}$ is the $P\times N$ matrix of stored patterns. This non-iterative solution makes the training process of KRR significantly faster than the gradient-based optimization required for KLR, a difference we quantify in our experiments.
@@ 3.4 Recall Process
The recall process in KLR-trained networks differs from that in Hebbian and LLR-based networks, as it does not rely on an explicit $N\times N$ weight matrix. Given a network state $\bm{s}(t)$ at time $t$, the next state $\bm{s}(t+1)$ is determined by the sign of each neuron’s activation potential relative to a threshold $\theta$:
@@ 3.5.1 Network and Learning Parameters
For the KLR learning process, we used gradient descent with a learning rate $\beta=0.1$ and number of parameter updates $M=200$. For the KRR learning process, the closed-form solution was used. The L2-regularization parameter $\lambda$ was set to a default value of $\lambda=0.01$ for most experiments, except for the sensitivity analysis. As shown in Appendix A, the learning curve for a challenging high-load condition ($N=500$, $P/N=4.0$) demonstrates that the loss substantially decreases within the initial $\sim$150 steps and approaches a plateau, confirming that 200 updates are adequate to achieve a well-converged model.
@@ 3.6.3 Quantitative Metrics
  • Fixed Point Rate: The proportion of trials that converged to any fixed point (including target, other learned, and spurious). This is equivalent to 1.0 - (Cycle Rate + Not Converged Rate).
@@ 4 Results
A foundational aspect of the network’s dynamics is its convergence behavior. While a formal proof of convergence for KLR-trained Hopfield networks with synchronous updates is not established, our experiments revealed a remarkable degree of stability. Across the tens of thousands of recall trials conducted for this study, every single trial converged to a stable state (either a fixed point or a limit cycle) within the maximum of 30 update steps.
Furthermore, a detailed analysis of non-stationary dynamics showed that the rate of convergence to limit cycles (Cycle Rate) was negligible ($0.1\%$) across all tested parameter regimes (see Appendix C for a representative plot). This empirical observation that the dynamics are overwhelmingly dominated by convergence to fixed points validates our subsequent focus on the analysis of these fixed point attractors and their properties.
@@ 4.1 Performance and Attractor Landscape Characteristics ($N=500$)
Figure 1: Performance and attractor landscape characteristics of the KLR-trained network ($N=500$). All curves show the mean values averaged across five master random seeds, with shaded regions representing the 95% confidence intervals. (a) Capacity Curve: Target Recall Rate as a function of storage load ($P/N$) for various initial similarities. The network maintains near-perfect recall up to $P/N=4.0$ for low-noise inputs. (b) Decomposition of Recall Errors: Recall Errors (1.0 - Target Recall Rate) are decomposed into convergence to other learned patterns and spurious attractors. Colors distinguish the error type, while line styles represent different initial similarities. For moderate noise levels, the rate of spurious errors is negligible, demonstrating a remarkably clean landscape where mis-recall to other learned patterns is the dominant failure mode. However, under extreme conditions (high load and high noise), spurious errors become significant. (c) Noise Robustness: Target Recall Rate as a function of initial similarity for various storage loads ($P/N$). At low loads, recall is robust even from highly corrupted states. (d) Convergence Speed: Average number of steps to convergence for various initial similarities. Recall is exceptionally fast, typically within 1-2 steps for high-similarity inputs and well under 10 steps even for noisy, high-load conditions.
@@
To understand the nature of the recall failures observed at high loads and noise levels, we decomposed the errors into their constituent sources. The results, shown in Fig. 1(b), reveal a crucial the detailed breakdown of recall failures.At low to medium storage loads ($P/N<1.5$), the rate of convergence to spurious attractors (Spurious Error, green lines) is negligible. In this regime, the dominant source of error is the convergence to an incorrect but valid learned pattern (Other Learned Error, orange lines). This “mis-recall” indicates that the basins of attraction for different stored patterns are competing with each other. However, a distinct behavior emerges at very high storage loads ($P/N>2.0$) under high noise conditions (initial similarity <0.2<0.2). As the storage load increases further, the rate of mis-recall to other learned patterns decreases, while the rate of convergence to spurious attractors sharply increases, eventually becoming the dominant failure mode. This suggests that in the extremely overloaded regime, the energy landscape becomes so rugged that the system frequently gets trapped in local minima (spurious states) that do not correspond to any stored memory, rather than being attracted to a wrong memory. It is important to note that for lower noise levels (e.g., initial similarity $\geq 0.4$), the Spurious Error remains low even at high loads, confirming the robustness of the model within a reasonable operating range.
@@
Finally, the recall dynamics are exceptionally fast, as illustrated in Fig. 1(d). For high-similarity initial states, the network consistently converges in just one to two steps. Even for highly corrupted states in the low-similarity regime, the average number of steps to convergence remains remarkably low, typically well under 10 steps. Interestingly, for these noisy inputs, the convergence time exhibits a non-monotonic “hump” shape, peaking at intermediate storage loads. The subsequent decrease in convergence time at very high loads corresponds precisely to the regime where recall failures dominate (cf. Fig. 1(b)). This suggests that the dynamics quickly settle into nearby stable states—either incorrect learned patterns or spurious fixed points—rather than undertaking a prolonged search for the correct target. This rapid convergence is a significant practical advantage of the model.
@@ 4.3 Scaling Analysis of Storage Capacity and Dynamics
Figure 3: Scaling analysis of storage capacity and convergence dynamics under N-dependent optimal kernel widths ($\gamma_{\text{opt}}$). The analysis was performed under a moderate noise condition (Initial Similarity = 0.8) for three different network sizes: $N=100,250$, and $500$. (a) Storage Capacity Scaling: The Target Recall Rate curves for all three network sizes largely collapse onto a single curve, particularly in the high-performance regime ($P/N\lesssim 3.5$). While minor deviations are observed at the highest loads, this trend provides strong evidence that the storage capacity $P$ scales linearly with the network size $N$. (b) Convergence Speed Scaling: In contrast, the convergence dynamics exhibit a clear dependence on the absolute network size. For a given $P/N$ ratio in the high-load regime, larger networks converge significantly faster. This suggests that while the stability of attractors is governed by relative load, the efficiency of the recall dynamics is enhanced in higher-dimensional state spaces.
@@
During our initial investigation, a naive application of the conventional kernel scaling rule, $\gamma=1/N$, led to the counter-intuitive result that the storage capacity coefficient $P/N$ decreased for larger networks. This motivated a focused analysis of the interplay between $N$ and $\gamma$. We found that the optimal kernel width, $\gamma_{\text{opt}}$, which maximizes recall performance at high loads, exhibits a non-trivial dependence on the network size. Specifically, our empirical analysis revealed that larger networks require a more localized kernel to achieve maximum capacity, with the optimal scaling factor $c_{\text{opt}}=\gamma_{\text{opt}}N$ increasing with $N$ (e.g., $c_{\text{opt}}\approx 2$ for $N\leq 250$ and $c_{\text{opt}}\approx 5$ for $N=500$). A detailed account of this hyperparameter optimization is provided in Section 4.4.
@@
As shown in Fig. 3(a), the Target Recall Rate curves for all three network sizes now largely collapse onto a single curve, particularly in the high-performance regime. While minor deviations are observed at very high loads (e.g., $P/N=4.0$), the performance is predominantly determined by the relative storage load $P/N$, irrespective of the absolute network size $N$. This result provides strong evidence that, under appropriately scaled kernel parameters, the storage capacity of the KLR-trained network follows the fundamental linear scaling law, $P\propto N$. This confirms that the high capacity coefficients reported are a general property of the model and not a finite-size effect.
@@ 5 Discussion
Our extensive quantitative analysis has provided a comprehensive picture of the attractor landscape in KLR-trained Hopfield networks. A central finding is the emergence of a remarkably “clean” landscape, characterized by a near-zero rate of spurious fixed points, particularly under high storage loads with moderate noise (Fig. 1(b)). This strong suppression of spurious states is a key factor behind the model’s high fidelity. But what is the mechanism behind this suppression?
@@ 6 Conclusion
Our findings confirm that KLR training sculpts a remarkably clean and robust attractor landscape, characterized by a near-zero rate of spurious states across a broad range of parameters and exceptionally fast recall dynamics. Our comparative analysis revealed that this high performance is a general feature of kernel regression methods, though KLR and KRR present a trade-off between iterative flexibility and closed-form efficiency.
@@
In summary, this work provides a clear set of empirical findings and practical design principles for constructing high-capacity, kernel-based Hopfield networks. By systematically characterizing the properties of the attractor landscape and its dependence on key parameters, we have laid a solid foundation for future theoretical work and explorations into alternative loss functions, such as hinge loss for sparsity-induced efficiency. This establishes a robust framework for the engineering of next-generation, brain-inspired memory systems.